wiki:UC20191202

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 12/02/19

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Kevin Schlaufman (JHU), Derek Buzasi (FGCU), Sarah Tuttle (UW), Eric Bellm (UW), Ben Williams (UW), John Bally (CU), Misty Bentz (GSU), Moire Prescott (NMSU), Mukremin Kilic (OU), Anne Verbiscer (UVa)


User feedback and comments from institutional representatives:

  • FGCU - nothing to report except being plagued by bad weather
  • Wyoming - no report
  • Oklahoma - nothing to report
  • NAPG - no report
  • Seattle - no report
  • Georgia State - nothing to report
  • UVa - nothing to report
  • Colorado - nothing to report
  • JHU - nothing to report
  • NMSU - nothing to report
  • Washington - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 10/29/2019 - 11/29/2019

0) Overview

We had minimal problems this period. The weather started out good with good observing efficiency and then later in the month we lost substantial time due to weather. We had several class visits this month and the annual BoG meeting in Chicago. We have made substantial progress in planning for a public open house on March 7 with a half-night of eyepiece viewing. It’s been a busy month with classes, traveling, KOSMOS II review, people out due to medical issues, and weather.

1) Telescope

Nominal operations with minimal problems. Some tracking issues remained but reduced in frequency the as seasonal temperature changes stabilized. Bill and Shane made progress in understanding Fritz’s axis controller system.

2) Instruments

DIS: The blue camera contamination is still bad and there are no plans to service it. The red camera is still getting worse slowly but is not considered bad yet. We did have a mounting issue with the upper part of the support cart on the telescope that we were able to improve substantially.

TripleSpec is operational. A pattern noise was noted at a low level on some images and UVa is looking into it. Ed found an issue with the connectors and that improved things quite a bit. The noise level is reduced but it still may be present at times. We are back at the monitoring stage right now.

Agile is operational.

Echelle is operational. The inter-order light In blue has improved. Another odd readout error occurred and we are monitoring it. Ecam windowing with binning is in work but at lower priority as it is a seldom used configuration.

NIC-FPS is operational.

ARCTIC is operational. We had one minor problem with an umbilical hose and also had to reset a lost quad once.


Additional telescope and instrument discussion:

Mark highlighted a few items that were touched on in his report. We've had more bad weather at the site recently, although we got back on sky last night. We survived a major winter storm fairly well, during which the biggest problems were wind (not snow) and freezing ice. For those of you who are new to APO, the main winter concerns for the 3.5m are a) at cold temperatures things (e.g. electronics, motors) don’t work well, b) snow and ice on the dome pose a threat to the telescope if opened, c) snow on the lightning arrestors means that the structures are more susceptible to lightning strikes, and d) condensation or freezing on the primary mirror is obviously bad for the optics. The site staff looks at all of these things to determine open/close conditions, and sometimes skies are clear but we still can’t open. We recognize that this is a frustration for the scheduled observers, but we need to keep the safety of all people and equipment on site as our highest priorities.

We held the KOSMOS II pre-ship review and several items were identified for follow-up work.

The status of DIS is effectively unchanged, and we are trying to avoid servicing it during winter months.

Regarding the TripleSpec pattern noise, changing a connector did help. The noise source appears to be possibly two different noise sources: one is a 60 Hz pickup noise, and the other is directly aligned with rows and columns (about 3 DN above the background). The UVa team is examining possible causes.


Q4 3.5m scheduling:

Q4 is fully scheduled; there is no more OPEN or DD time available.


Q4 ARCSAT scheduling:

There is still some OPEN time remaining in December. Contact Russet, Ben, and Nancy if interested in using it.


Q1 3.5m scheduling:

We have received the Q1 requests from almost everyone (still waiting for one more), and Russet will be assembling the schedule this week. Update: the Q1 scheduled is published! Users interested in requesting any of the OPEN time should contact Ben, Nancy, and Russet and include a completed proposal form if there is not already one in the system for Q1.


Q1 ARCSAT scheduling:

The call for Q1 ARCSAT proposals will be issued once the Q1 3.5m schedule is completed. Update: the call for Q1 ARCSAT proposals is out, and there is still a lot of time available. Users interested in requesting any ARCSAT time in Q1 should contact Ben, Nancy, and Russet.


NA2 Baffle Project:

The NA2 baffle is now in fabrication and is expected to be delivered to APO in late January 2020. Update: the observing staff is developing a commissioning plan, which will include before-and-after observations made to assess vignetting and scattered light with both the old and new baffle.


KOSMOS Update:

The scope of the recent pre-ship review included everything outside of the dewar, with the goal of identifying any issues that should be resolved in the lab. The review was very helpful, with the UW Telescope Engineering Group providing feedback on questions that had arisen, and they accommodated the requests for some modifications needed to making handling the instrument easier for the obs-specs. Due to fact that the detector suffered some damage, the decision was made to postpone integration at APO. The detector has some broken wire bonds; it was sent to the detector lab in Tucson but they could not repair it. It has now been sent back to e2V (Teledyne), but they are moving quite slowly. For the time being the TEG will use the engineering detector in the lab to get it installed to check out communications, software, etc. Next week they will make an assessment on when the instrument might be shipped, and whether it will be with the engineering detector or not. Update: the TEG is working in parallel on three issues: resolving a vacuum leak with the new dewar, getting a status update re: our detector from e2V, and identifying backup options in the event that the science detector has to be replaced. Still no word on a ship date, and this will likely not be resolved until the dewar issue is fixed and some lab tests are done with the engineering grade detector.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: No update. Nancy will request a status update prior to the next telecon.
  • Getting narrow-band filters for ARCTIC: Nancy will make a decision so she can get them ordered before the end of CY19. Update: these have not yet been ordered, but we are requesting quotes for a larger format, narrower H-alpha filter and an [OIII] 5007A filter.
  • Science communications intern: Nancy hired two NMSU undergraduates, Hannah Perry and Emily Medeles, for this position. Hannah is a senior majoring in journalism and mass communications with a minor in advertising, and Emily is a sophomore Physics major and Astronomy minor. Together this dynamic team will be helping us to boost the social media presence of the 3.5m, develop new marketing materials, and promote science highlights resulting from 3.5m observations. You will be hearing more from them in the coming months!
  • AAS presence: Nancy reserved an exhibitor booth, and needs to know who from ARC and leasing institutions is a) going to the AAS meeting in January and b) willing to help out by spending some time at the booth talking about APO to passers-by. After polling the UC reps on the call and based on other input Nancy received, the following people are going and willing to spend time at the booth: Nancy, Ben, Mike Crenshaw, Karen Kinemuchi, Nickalas Reynolds. Update: the booth was a success! Stay tuned for more info at the February Users Committee telecon.
  • Student training videos: Nancy is in discussions with some of the observing specialists and Shane regarding what kinds of training videos can be made in-house. No further updates concerning asking class visitors to film themselves learning a specific task.

ACTION ITEMS:

Open action items from previous meetings:

  • Nancy will contact trip leaders bringing student groups to APO in Q4 to discuss training video ideas. STATUS: OPEN.
  • UC reps should poll their users to see who is going to the AAS meeting and would be willing to help out with booth presence. STATUS: CLOSED.
  • UC reps should poll their users, particularly those with a vested interest in an echelle replacement, and ask for volunteers to serve on the echelle replacement advisory group. STATUS: CLOSED

New action items from this meeting:

None.


All Other Business

None.


Next meeting:

The next Users Committee telecon would normally be on January 6, but that is during the AAS meeting. Nancy will send an email to schedule the Spring 2020 calls. We will try using Zoom for these meetings to see how we like it. Nancy will send out the Zoom meeting invitation before each call. Update: we will alternate between the first Tuesdays and Mondays of the month between February and June 2020. Thus, the next Users Committee call will be on Tuesday Feb. 4 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.